. Find the square of the following numbers. i. 32  
ftjstrfjmtrjkykmyfkjrtfjrsdhgedgbedzhbdeh

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

i. (32)2
= (30 +2)2
= (30)2+ (2)2+ 2×30×2 =900+4+120

= 1024



Thanks